# Cole: A Hierarchical Generation Framework for Graphic Design

### 28 Nov 2023 v1

@@ C. The Detailed List of DesignerIntention
A comprehensive list of designer intention prompts is summarized across Table 4, Table 4, Table 4, and Table 6.
@@
Table 4 details the intention prompts corresponding to the images exhibited in Figure 1 of the main paper. In Figure 20, we show a few example pairs that illustrate the original image information alongside the corresponding generated intention prompt.
@@
Table 4 presents the intention prompts augmented with GPT-, which were utilized for DALLE3 in Figure 2 of the main paper.
@@
In Table 4, we display the augmented intention prompts with GPT- that were employed for DeepFloyd/IF, SDXL, and DALLE3 in Figure 10 of the main paper.
Last, Table 6 provides the detailed intention prompts for the design images generated, as depicted in Figure 11 of the main paper.
D. More Ablation Experiments
@@
| Category |
Intention Prompt |
| marketing | Create a poster advertising a corner spice shop with a wide range of spices and ingredients. |
| posts | Promote a stand-up comedy show happening on November 12th, 20xx. |
| marketing | Design a gift certificate for a travel insurance company promoting their services with a 30% discount offer. The certificate should convey a sense of safety and protection while travelling. |
| posts | Design a recruiting post for a graphic designer job opening. Encourage interested candidates to send their CV to yourname@email.com |
| marketing | Design a Christmas sale poster with a fun and festive look. The poster should advertise up to 50% off on Christmas accessories, with a limited time offer. |
| posts | Create a LinkedIn post with a cartoon illustration to promote an article about the benefits and growing popularity of remote work from home. The post links to a website for more information about the topic and is associated with a specific company and CEO John Doe. |
| marketing | Create a poster advertising babysitting services. The design should be warm and inviting, featuring a little girl and emphasizing professional and reliable childcare. The business name and phone number should be prominently displayed (+123 456 7890). |
| marketing | Create a poster advertisement for a restaurant’s happy hour special, showcasing the luxury and elegance of the establishment. The poster should include the restaurant’s name, location, website, and phone number for reservations and delivery. Happy Hour runs from 7 PM to 11 PM. |
| marketing | Create a real estate rental property flyer with a cozy and inviting interior design. Showcase the apartment’s features, amenities, and price. Include contact information for the real estate agency and their website address. The monthly rental price is $13,950. Bedrooms: 2, Bathrooms: 3, Square Footage: 1500. |
| creative | Design an interactive digital invitation for a space-themed gala. The invitation should feature an animated rocket launch sequence and include a scannable QR code that plays a personalized greeting from a holographic astronaut. |
| creative | Create a series of posters for a fictional Ancient Greek Olympics, with each poster highlighting a different sport like chariot racing or discus throw, styled in traditional Greek pottery art. |
@@
| Category |
GPT4 Refined Intention Prompt |
| events | A vibrant birthday card bathed in festive hues of red, orange, and yellow. A delectable cake, crowned with an assortment of juicy fruits, serves as the focal point. Around it, confetti and streamers tumble playfully. The card invites you to a birthday celebration on June 22nd, from 8pm to 11pm, in a quaint house located at 123 Street, Any City 4567. |
| marketing | Picture a vibrant gift certificate for a leading travel insurance company. A globe subtly watermarked in the background, planes, and ships creatively darting across. Emblazoned on the center in bold, gold lettering is the irresistible offer: ”30% Discount on All Travel Insurance Policies”. Borders of the certificate are adorned with famous landmarks, adding a touch of wanderlust. The company’s logo sits proudly at the top, promising security and peace of mind for all journeys. |
| posts | A vibrant autumnal landscape with leaves in shades of gold, amber, and crimson cascading from the trees. A sleek banner sweeps across the scene, displaying the text ”30% off Autumn Collection”. A stylish arrow gently points downwards towards a glowing ”Link in Bio” button. The whole image exudes an irresistible allure, beckoning Instagram viewers to partake in this exclusive offer. |
| covers& headers | A vintage-inspired thumbnail for a vlogger episode, prominently featuring a curated assortment of five rare antique telephones, each unique in design and era. The collection is artfully arranged on a mahogany table against a backdrop of antiquated maps and time-worn books. The eye-catching title, ’Top 5 Antique Phones’, is elegantly superimposed on the scene in a classic, bold typeface. |
| creative | Imagine a captivating poster for a blues concert nestled in a moonlit clearing of the Amazon Rainforest. The midnight blue canvas paints an enchanting scene of silhouetted blues instruments - a saxophone, a guitar, and a harmonica - intertwined with the silhouettes of nocturnal wildlife like sleek panthers and hooting owls. The artwork pulsates with the soulful rhythm of blues music, echoing through the wilderness, stirring the soul. |
@@
| covers& headers | Create a Youtube thumbnail for a video showcasing how to make Italian dishes at home, specifically focusing on rolling out dough. |
| covers& headers | Create a Youtube thumbnail to promote a pastry making tutorial series. The series is called ’The Best Pastry in the World’ and is hosted on the channel ’Your Channel’. The thumbnail should showcase a collage of homemade pastries, specifically croissants. Part 01 should be emphasized. |
@@
| covers& headers | Design a Facebook event cover for a winter decoration workshop featuring lectures and demonstrations by Miranda Yates, Sam Knightly, Zachary Evans, and Veronica Floyd on December 24 at Lafayette Studio. |
@@
| marketing | Create a Thanksgiving gift certificate with an elegant and sophisticated design that offers a 50% discount on decorations. The message wishes a happy Thanksgiving to the recipient. |
| marketing | Create a coupon that promotes a garden design service offered by Yard Culture. The coupon is a gift voucher valued at $49 and can only be used once. It is valid until November 15th. The coupon number is No:7496-4587. |
| marketing | Design a poster to promote a charity bike ride event to raise funds and support for a humanitarian cause. The event will be held on June 26th and the website for registration and donation is www.funbike.com. |
@@
| posts | Create a recruiting Instagram post to advertise a job vacancy for a UI/UX designer. Encourage passionate individuals to apply and provide a link to apply in the bio. |
| posts | Create an Instagram story promoting a special Christmas offer for a chocolate drink. The offer includes a 20% discount and encourages customers to order now. |
@@
| events | Design a holiday-themed gift card to promote the sale of gift boxes and packaging for the Christmas in July event. Include the website address www.yourbrand.com and store name. |
@@
| events | Create a business card for a car service company that highlights their commitment to providing reliable and efficient car maintenance and repair services. The slogan ’Give your car wings’ emphasizes the quality of service provided. |
| events | Design an advertisement card to promote the sale of premium quality school equipment, with a focus on enhancing the learning experience for students. |
| events | Create a Women’s Day greeting card with signs of female gender, celebrating women and their achievements. The design should be beautiful and include floral elements. The card should be suitable for store use and have a message of ’Happy Women’s Day’. |
| advertising | Create an Instagram ad for a real estate agency highlighting the modern and prestigious skyscraper building they have on offer. Encourage viewers to ’reach new highs’ by visiting the website at www.togalim.eu. Include contact phone numbers 126-3764-1376 and 301-17361-5085 for interested clients. |
| advertising | Create an Instagram ad for a real estate agency highlighting the modern and prestigious skyscraper building they have on offer. Encourage viewers to ’reach new highs’ by visiting the website at www.togalim.eu. Include contact phone numbers 126-3764-1376 and 301-17361-5085 for interested clients. |
| advertising | Create a Facebook advertisement for a store selling natural essential oils that promote healing for the body and soul. Highlight the use of floral and plant ingredients in the product. Encourage viewers to visit the store. |
| advertising | Design a weekend sale advertisement for skincare products with a focus on promoting the cosmetic serum. All items are discounted up to 30% for this season. |
| advertising | Create an Instagram ad for a camping festival and hiking event. The ad should promote the event and provide information about registration, including a website. Additionally, offer tips for surviving while hiking. |
| advertising | Create an Instagram ad for a toy store’s Black Friday and Cyber Monday sale with a focus on the discount of 45% off on selected toys. |
| creative | Design an attractive poster for a Mars botanical garden tour. Illustrate a greenhouse with a variety of extraterrestrial plants. Use vibrant colors to depict the potential biodiversity on Mars. |
| creative | Create an enticing poster for a classical orchestra concert in the heart of the Amazon Rainforest. Depict the harmony between the music and the wilderness with a blend of vibrant greens and subtle sepia tones. Incorporate subtle hints of wildlife and flora within the musical instruments. |
| creative | Design a dramatic poster for an opera performance in a natural Amazon Rainforest amphitheater. Use a bold, Baroque-inspired style to reflect the grandeur of opera, juxtaposed with the raw beauty of the forest. Show a dramatic scene unfolding on the forest stage. |
| creative | Create a poster for an imaginary music festival where classical composers like Mozart and Beethoven are headliners. The design should merge classical elements with a modern music festival vibe, including stage setups and audience. |
| creative | Construct an eye-catching poster for an upcoming academic conference on Quantum Physics. The design should encapsulate the mysterious and profound nature of the subject. Use a black hole or quantum particles as the main design element, and apply a deep space color scheme to create an intriguing atmosphere. Make sure the conference details and keynote speakers are clearly visible. |
| Category |
Image Generate Prompt |
| marketing | Gift Certificate. blank, empty, layout, tb, certificate, offer, illustration, pattern, floral, flowers, pink, frame, sale. The image features a white background with a pink border. The border is adorned with a decorative floral pattern, giving the image a visually appealing and elegant appearance. The combination of the white background and the pink border with the floral pattern creates a harmonious and attractive design. |
| posts | Instagram Story. astronaut, cosmic, cosmonaut, helmet, orange, science, space, spaceman, universe, story, instagram, ig, insta, frame. The image features a man wearing a spacesuit, sitting on a planet with his legs crossed. He is positioned in the center of the scene, and the planet is displayed in the background. The man appears to be a spaceman, possibly an astronaut, as he is dressed in a suit designed for space travel. The scene is set against a dark blue sky, which adds to the overall atmosphere of the image. |
| marketing | Gift Certificate. agency, holiday, journey, service, tourism, travel, traveling, trip, vacation, tb, certificate, sale, discount, insurance, illustration, cartoon, blue. The image features a map with a variety of travel-related items scattered around it. There are several pairs of shoes, including a couple of boots, placed in different positions on the map. A hat is also present, positioned near the top left corner of the map. In addition to the shoes and hat, there are multiple books scattered across the map, with some located near the bottom left corner and others in the middle and top right areas. A pair of scissors can be seen in the top right corner of the map, and a clock is placed near the center of the map. |
| events | Business Card European. data, card, health, home, professional, service, technology, icon, house, nursing, care, modern, medical, health care, contact, information. The image features a house with a pink roof, which is situated next to a street. The house is surrounded by a white fence, and there is a person standing on the sidewalk in front of the house. The person appears to be looking at the house, possibly admiring its architecture or considering it for purchase. The scene also includes a car parked on the street, adding to the overall atmosphere of the neighborhood. |
| events | Instagram Post. wedding, dinner, rehearsal, banquet, event, holiday, celebration, marriage, food, party, restaurant, ceremony, leaf, jungle, tropical, exotic, brown, instagram, post, insta, ig, tb. The image features a brown background with a few green leaves scattered throughout the scene. The leaves are of various sizes and are positioned at different angles, creating a sense of depth and texture. The overall composition of the image is simple, with the brown background and green leaves being the main focus. |
| events | Instagram Post. empowerment, equality, feminism, holiday, march, march 8, rights, women, women’s day, womensday, tb, illustration, female, girls, celebration, beautiful, greeting, sisterhood, instagram, insta, ig, post, flowers, floral. The image features two people standing next to each other, each holding a heart. The hearts are positioned above their heads, creating a visually striking scene. The people are standing in front of a backdrop of flowers, adding a touch of color and beauty to the image. |
| covers& headers | Ticket DIN large. festival, art, abstract, geometric, graphic, illustration, pattern, shape, simple, visual, colorful, blue, ticket. The image features a colorful and abstract design, with a predominantly blue background. The design consists of various geometric shapes, including triangles, squares, and circles, arranged in a visually appealing manner. The shapes are positioned throughout the image, with some closer to the foreground and others further in the background. The overall composition creates a sense of depth and complexity, making the image visually engaging and interesting. |
| advertising | Square Graphic Post. tb, instagram, insta, ig, post, style, fashion, trendy, fashionable, stylish, beautiful, clothes, model, woman, posing, modern, female, young, casual, beauty, attractive, look, color, modeling, lady, people, outfit, order, new arrival, pink, customer, testimonial, review. The image features a pink background with a white square placed in the center. The square is surrounded by a pink hue, giving it a slightly pinkish tint. The background is predominantly pink, with no other objects or distractions visible in the scene. |

### 18 Mar 2024 v2

@@ C The Detailed Prompt List of DesignerIntention
A comprehensive list of designer intention prompts is summarized across Table 6, Table 6, Table 7 and Table 8.
@@
Table 6 details the GPT-4 augmented intention prompts corresponding to the images exhibited in Figure 1 of the main paper.
@@
Table 6 presents the intention prompts which were utilized for Cole in Figure 2 of the main paper.
@@
Table 7 presents the intention prompts and GPT-4 augmented intension prompts which were utilized for Cole in Figure 11 of the main paper.
Last, Table 8 presents the intention prompts which were utilized for Cole in Figure 12 of the main paper.
D More Details of Bin Partition Scheme within Typography-LMM
To streamline the Typography-LMM’s prediction process, we adopt a discretization strategy, i.e., bin partition scheme, for continuous floating-point values based on the approach described in [6]. This involves categorizing continuous variables into discrete bins and representing the values within each bin with a unique token. The specifics of our bin partitioning scheme are detailed below:
Font Size: The range [2, 200] is divided into 100 sub-intervals.
Angle: The range [0, 2] is segmented into 64 sub-intervals.
Color (RGB values): The range [0, 255] is split into 32 sub-intervals.
Text Bounding Box Coordinates: The normalized coordinates within the range [-1, 1] are partitioned into 256 subspaces.
Opacity: The range [0, 255] is divided into 8 subspaces.
Letter/Line Spacing: The normalized values within the range [0, 1] are divided into 40 sub-intervals, adjusted according to the image width.
We estimate the start and end points of all the above partitions based on the statistics derived from the training set, accommodating a broad spectrum of typography characteristics, including font size, angle, letter and line spacing, opacity, dimensions of the font bounding box (left, top, width, height), and color.
E More Ablation Experiments
Design LLM We calculate the semantic scores (semantic textual similarity scores estimated based on Sentence-BERT) to compare the ground-truth with predictions from our Design-LLM, GPT-3.5, and GPT-4 across the test dataset. To ensure accurate predictions, we supply three in-context examples for both GPT-3.5 and GPT-4. Furthermore, we restrict GPT-3.5 and GPT-4 to generate values for only one key query at a time, simplifying the task’s complexity. The results, detailed in Table 4, reveal that our Design-LLM significantly outperforms both GPT-3.5 and GPT-4.
| Keys in JSON | Design LLM | GPT-3.5 | GPT-4 |
| Global Caption | 68.57 | 56.31 | 59.56 |
| Category | 80.32 | 80.93 | 78.65 |
| Keywords | 89.31 | 84.67 | 84.14 |
| Background Caption | 57.69 | 48.06 | 52.37 |
| Object Caption | 73.59 | 73.21 | 73.16 |
| Text layer (heading, sub-heading, body-text) | 54.75 | 46.80 | 53.56 |
@@ E More Ablation Experiments
| marketing | An elegantly designed gift certificate for a reputable travel insurance company. The certificate, made of high-quality paper, is adorned with motifs of different modes of travel - airplanes, ships, and trains. A shiny, golden 30% off stamp is prominently placed at the center. The company’s logo is tastefully positioned at the top, while the certificate’s edges are softly curved, enhancing its appeal. |
| covers& headers | A vividly illuminated stage under a starry night sky, acting as the primary backdrop. Smack in the center is a caricature of John Oliver, his signature glasses glinting, holding a microphone with a wide grin. At the top, bold letters spell ”Laugh Out Loud with John Oliver”. Below, in elegant script, is the location ”At The Comedy Club, New York”. A chuckling audience silhouette completes the lively atmosphere. |
| creative | A collection of vibrant posters for a fictitious Ancient Greek Olympics. Each poster features a different sport: a fiercely competitive chariot race with horses kicking up dust, a discus thrower showcasing his strength against a setting sun, all stylized like traditional Greek pottery art, with the sportsmen clad in traditional tunics, their muscles rippling under the warm, orange glow of the Grecian sun. The borders of each poster are adorned with intricate patterns of laurel wreaths. |
| advertising | An eye-catching weekend sale advertisement for skincare products, featuring prominently a sleek bottle of cosmetic serum bathed in a soft, ethereal glow. A bold, ruby-red text proudly declares ”All Items Discounted Up To 30%”. Surrounding the serum are various skincare items, all discounted, framed against a backdrop of vibrant autumn leaves, reflecting this season’s promotion. |
@@
| Category |
Intension Prompt |
GPT-4 Augmented Prompt |
| advertising | Create a customer testimonial advertisement showcasing a pink clothing collection. The advertisement includes a positive review from Jenny Wilson. | A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged. In the corner, a gleaming starburst accentuates a glowing testimonial from Jenny Wilson, her words of praise shimmering. Beneath the testimonial, elegantly scripted contact information is highlighted: phone number 123-456-789, website WWW.FASHIONSTYLE.COM, and an address at 123 Anywhere ST., Any City. |
| events | Create a birthday card with a festive and colorful design featuring a cake with fruits. The card is an invitation to a birthday party happening on June 22nd from 8pm to 11pm at 123 st. Any city 4567. | A burst of vibrant pink hues dominates the canvas: a mix of blush, fuchsia and rose clothing neatly arranged. In the corner, a gleaming starburst accentuates a glowing testimonial from Jenny Wilson, her words of praise shimmering. Beneath the testimonial, elegantly scripted contact information is highlighted: phone number 123-456-789, website WWW.FASHIONSTYLE.COM, and an address at 123 Anywhere ST., Any City. |
| creative | Design an alluring poster for a blues concert in a clearing of the Amazon Rainforest. The artwork should evoke the soulful tunes of blues music echoing through the wilderness. Use a moody, midnight blue palette and incorporate silhouettes of blues instruments and nocturnal wildlife. | An enticing poster for a blues concert set in an Amazon Rainforest clearing. The artwork is bathed in a mysterious, midnight blue hue, resonating the soulful melody of blues music. The scene is punctuated by silhouettes of blues instruments - a haunting saxophone, a melancholic guitar, and a poignant piano. The wilderness comes alive in the shadows of nocturnal creatures - a stealthy jaguar, an elusive tree frog, and a silent owl perched on a branch, all harmoniously intertwined with the rhythm of the blues. |
| Category |
Intention Prompt |
| posts | Design a funny and creative Halloween Facebook post with humorous messages and stickers. Promote a Halloween party at Under Bar with the caption ’Get your treat, Trick or Treat, dude!’ |
@@
| creative | Create a recruitment post for a company seeking an illustrator to join their team. |
| posts | Create a Facebook post featuring a happy family with their child enjoying a leisurely day outdoors. Capture their stylish outfits and cute smile. Frame the image with a white border and add a shadow effect. |
| posts | Create a Facebook post to promote a special menu for pancakes at a cafe or restaurant. The offer is delicious pancakes for only $10. The website for more information is www.pancakehouse.com. |
| posts | Create a happy and joyful image of a pregnant woman enjoying a relaxing day by the seacoast. The text emphasizes the excitement and love for the baby growing inside. |
@@
| posts | Design a Facebook post promoting a summer collection for women. Highlight the new arrival and offer a 60% discount for this week only. Encourage viewers to shop now. |
| posts | Design a Facebook post advertising an organic cosmetics kit with natural and organic ingredients. The kit is a must-have for skincare and wellness. |
| posts | Promote a cheesy Italian pizza with a discount offer of 30% for the first order. Encourage customers to visit the restaurant or order online. The promotion is for a limited time only. |
| covers& headers | Design an advertisement for a crowdfunding platform called LocalBless. Encourage users to contribute to social projects by visiting the website and donating. Create a visually appealing image with a stone pattern background and triangles to convey a sense of stability and trust. |
| posts | Design a Facebook post promoting a two-week summer camp that provides hands-on, project-based instruction in tech. The camp will take place from 12-14 August. |
| posts | Create a minimalistic and modern home design discount offer for a website or social media post. Encourage users to visit the website www.dreamhome.com to enjoy a 10% discount on their first design. |
| covers& headers | Create an attractive YouTube thumbnail for a makeup review video featuring an African American woman. The video will showcase the woman’s favorite makeup products and provide a weekly review. |
@@
| covers& headers | Create a Youtube thumbnail for a video about real estate investments, providing tips from professionals. |
| posts | Create a Facebook post promoting USA Independence Day tours with a patriotic theme. Encourage people to find their freedom in life by traveling more. Mention the offer of free cancellation. Include the website patrioutus.com for more information. |
@@
| posts | Create a Facebook post to celebrate International Pet Day with a cute little girl and a fluffy rabbit. |